# MNSC-0001-2024.00

### 15 Mar 2026 v1

@@
[…]
MNSC-0001-2024.00
[…]
Ahn and Jiang
[…]
Strategic Early Liquidation Strategies
[…]
Robust Optimal Strategies for Early Liquidation
in Financial Systems
[…]
@@
[…]
Systemic risk, Fire sales, Strategic liquidation, Worst-case approach, Permanent price impact, Interbank exposures
1 Introduction
The global financial crisis in 2007-2008 provoked intense discussions on the factors that threatened the stability of the financial system. In the majority of relevant studies, two channels of risk contagion have been widely acknowledged as significant factors contributing to this instability. One such channel involves direct debt exposures among financial institutions. When an institution fails to meet debt obligations due to adverse economic shocks, its creditor institutions experience financial distress due to the resulting repayment shortfall (EisenbergNoe2001). Furthermore, during the crisis of asset erosion, financial institutions with inadequate liquid assets are forced to sell off substantial illiquid assets to satisfy their debt claims. Such liquidation overwhelms shallow market demand and triggers sharp declines in asset prices. This phenomenon, termed price impact, imposes significant losses on institutions holding the same illiquid assets, even without direct debt exposures among them. These fire sales constitute another channel of risk transmission and are recognized as a salient ingredient of the crisis (ClercEtAl2016). Although the first channel has been studied more extensively in the literature, the second channel has gained increasing attention in recent years, particularly as financial institutions have actively reduced their debt exposures to other banks since the said crisis.
Our study departs from a conventional assumption in the existing literature that liquidation is conducted only at clearing. This assumption stems from the observation that liquidity-constrained institutions tend to favor borrowing over asset sales to avoid the prohibitive costs of liquidation. However, since borrowing capacity is not always guaranteed, relying solely on this option exposes banks to future funding shortages, particularly during systemic crises when credit conditions deteriorate (AcharyaEtAl2011). In this case, preemptive liquidation prior to clearing becomes a rational solution. Then, this naturally raises a critical question: what volume of illiquid assets should an institution sell early to ensure its solvency? The answer is not immediately evident as liquidation outcomes are driven by the interplay of institutional liquidation decisions that jointly determine sale prices and are constrained by repayment requirements.
@@ 1 Introduction
2 Literature Review
Impact of fire sales on systemic risk.
@@ 2 Literature Review
Our research builds upon the literature on risk contagion via fire sales of illiquid assets, initiated by CifuentesEtAl2005. They investigate the equilibrium of liquidation prices in financial systems where banks are forced to liquidate assets to satisfy regulatory constraints. ChenEtAl2016 formulate this equilibrium as an optimization problem and introduce a “liquidity amplifier” via sensitivity analysis, which quantifies how illiquidity intensifies fire-sale effects, thereby offering policy implications for intervention. In addition, AminiEtAl2016c establish sufficient conditions for the uniqueness of the equilibrium liquidation price, while WeberWeske2017 characterize the liquidation price in more complex settings where bankruptcy costs and cross-holdings are taken into account. These papers mainly focus on the assessment of systemic risk rather than the decision-making in liquidation and assume that liquidation takes place only at clearing. By contrast, we accommodate liquidation that occurs before clearing and offer novel insights into strategic liquidation decisions. A distinct line of work, pioneered by GreenwoodEtAl2015, leverages correlations in banks’ asset exposures to measure fire-sale spillover effects. Subsequent studies have enriched this paradigm by modeling the dynamic process of fire sales (DuarteEisenbach2021) and incorporating partial information about asset holdings (PangVeraart2023). Although their approaches open up new directions, our modeling focus remains on extending the framework of CifuentesEtAl2005, leaving the other models for further research.
Strategic liquidation in financial systems.
@@
Our work contributes to the growing literature of strategic liquidation under systemic risk. Feinstein2017 studies a financial system with multiple illiquid assets in which firms strategically choose asset sales to maximize their own valuations, and establishes the existence of equilibrium liquidation strategies. To analyze risk contagion arising from strategic liquidation, BraouezecWagalath2019 consider a static game in which banks maximize their expected profits subject to regulatory capital requirements and characterize the resulting Nash equilibrium. BichuchFeinstein2019 formulate an optimization problem in which borrowing and liquidation decisions are jointly determined to minimize total financing costs. While these studies analyze strategic liquidation through ex-post equilibrium outcomes of institutional interactions assuming rational behaviors of institutions, our work focuses on the ex-ante liquidation decision of an individual institution adopting a worst-case approach to the characterization of early liquidation strategies. [^1] It is essential to distinguish the notion of strategic liquidation considered here from the classical literature on optimal execution (e.g., BertsimasLo1998, GatheralSchied2013), which focuses on minimizing trading costs when unwinding a fixed asset position over a given period under exogenous market illiquidity. In our systemic context, by contrast, illiquidity is not merely a frictional cost but an endogenous channel of risk contagion that amplifies price declines and threatens systemic stability.
The role of permanent price impact in financial systems.
@@
Our research is closely related to studies on liquidation-induced systemic risk in the presence of permanent price impact, where price declines caused by liquidation do not immediately reverse and losses from liquidation accumulate over time. ContSchaanning2019 incorporate such cumulative impacts to empirically measure indirect contagion from fire sales empirically. CapponiEtAl2020b examine a dynamic setting where persistent price impacts can trigger a self-reinforcing cycle of mutual fund redemptions and further liquidations. In the context of collateralized systems, GhamamiEtAl2022 and PangVeraart2025 model clearing processes in which collateral is liquidated in multiple rounds, with price impacts from early sales carrying over entirely to subsequent rounds. While these works assume fully permanent price impact, implying that depressed prices do not recover, our work allows for partial price recovery. Furthermore, we examine robust liquidation decisions under permanent price impact from an individual decision-making perspective, whereas the aforementioned studies primarily focus on measuring systemic risk or elucidating contagion mechanisms.
Contagion through interbank exposures.
@@
A large body of literature on systemic risk investigates the impact of interbank obligations. The seminal framework of EisenbergNoe2001 models interbank repayments using a fixed-point characterization. Subsequent research has extended this framework along several dimensions while retaining its core analytical structure. For instance, RogersVeraart2013 incorporate bankruptcy costs into the framework and explore incentives for consortium-based bailouts. AcemogluEtAl2015 and CapponiEtAl2016 investigate how network topology and concentration affect shock transmission, respectively. KusnetsovMariaVeraart2019 and Veraart2020 extend the framework to accommodate multiple maturities and quantify distress versus default contagion, respectively. Furthermore, this approach has been applied to diverse areas such as systemic stress scenario selection (AhnEtAl2023), debt valuation (BaruccaEtAl2020, BanerjeeFeinstein2022), intervention policies (Ahn2019, CapponiBernard2022), and multilateral netting effects (Ahn2020, AminiEtAl2020). We also incorporate interbank exposures as an additional risk contagion channel. However, in contrast to the aforementioned works that analyze systemic risk from a regulatory perspective, our work examines the impact of this contagion channel on early liquidation decisions from the standpoint of an individual bank.
3 Research Questions
@@
An illustrative example.
Consider a system of $m$ banks operating over two periods ($t=1,2$), where liabilities are cleared at the end of the time horizon. Each bank $i$ possesses liquid assets valued at $e_{i}$ and holds $z_{i}$ units of illiquid assets. If the liquid assets prove insufficient to meet liabilities $L_{i}$ (i.e., $e_{i}<L_{i}$), the bank must cover the shortfall by selling off its illiquid holdings at a depressed price—a process known as fire sales. In contrast to the standard literature that assumes no early fire sales, we allow each bank $i$ to strategically liquidate its illiquid holdings $x_{i}$ at $t=1$, prior to clearing at $t=2$, in order to mitigate price impacts at clearing and sell a smaller amount of illiquid assets than in the scenario without early liquidation. Given a realization ${\bf x}=(x_{1},\ldots,x_{m})^{\top}$ of all banks’ early liquidation decisions, let $\psi_{i}({\bf x})$ denote the amount of illiquid assets bank $i$ must sell at clearing to fulfill its debt obligation for each $i=1,\ldots,m$. Then, by defining ${\bf x}_{-i}\coloneqq(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{m})^{\top}$, bank $i$’s liquid asset value at clearing becomes:
@@
Robust decision-making.
In this example, each bank $i$ would seek to determine its strategy $x_{i}$ that optimizes $\phi_{i}(x_{i},{\bf x}_{-i})$ while maintaining robustness against other banks’ strategies ${\bf x}_{-i}$, as these are practically unobservable. However, as well documented in the literature (AminiEtAl2016c) and as we shall demonstrate later, the mapping $\psi_{j}(\cdot)$ is characterized by a fixed-point equation involving nonconvex functions and is hence analytically intractable. A natural approach to this issue would be to replace $\psi_{j}({\bf x})$ by its surrogate $\psi_{j}({\bf 0})-x_{j}$ for all $j$, where $\bf 0$} denotes the zero vector. [^2] The early liquidation size $x_{j}$ is capped at $\psi_{j}({\bf 0})$ for all $j=1,\ldots,m$. This is because, as noted earlier, the goal of early liquidation is to reduce the amount of illiquid assets sold and the losses from liquidation compared to the case without it. Similarly, $\psi_{j}({\bf x})$ is capped at $\psi_{j}({\bf 0})-x_{j}$. This accounts for the worst-case scenario in which banks eventually sell off the same amount of illiquid assets as they would in the absence of early liquidation. Then, the said decision-making problem of bank $i$ can be formulated as follows:
@@
Solving (2) is surprisingly simple. By letting $s_{i}=\sum_{j\neq i}x_{j}$ and $\bar{s}_{i}=\sum_{j\neq i}\bar{x}_{j}$, it can be recast as $\max_{x_{i}\in[0,\bar{x}_{i}]}\min_{s_{i}\in[0,\bar{s}_{i}]}g_{i}(x_{i},s_{i})$, where $g_{i}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q(\bar{x}_{i}+\bar{s}_{i}-(x+s_{i}))$. Under the common assumptions on $Q(\cdot)$ in the literature (see Assumption 4.1), $g_{i}(x,s)$ is concave in $x$ and convex in $s$. Therefore, the minimax theorem (Rockafellar2015, Theorem 36.3) holds: $\max_{x_{i}\in[0,\bar{x}_{i}]}\min_{s_{i}\in[0,\bar{s}_{i}]}g_{i}(x_{i},s_{i})=\min_{s_{i}\in[0,\bar{s}_{i}]}\max_{x_{i}\in[0,\bar{x}_{i}]}g_{i}(x_{i},s_{i})$. Furthermore, $h_{i}$ satisfies the following point symmetry: $g_{i}(x_{i},s_{i})=g_{i}(\bar{x}_{i}-x_{i},\bar{s}_{i}-s_{i})$ for any $x_{i}\in[0,\bar{x}_{i}]$} and $s_{i}\in[0,\bar{s}_{i}]$}. Accordingly, $(\bar{x}_{i}/2,\bar{s}_{i}/2)$ is the saddle point of $h_{i}$, implying that
@@
Main issues and questions.
The simplicity of identifying the robust optimal strategy, however, is limited to the special case presented above. When practical considerations are incorporated, finding the strategy becomes significantly more complex. Specifically, this paper focuses on two distinct extensions prevalent in the literature: incorporating permanent price impact and including interbank exposures. In the former case, early liquidation exerts downward pressure on the asset price at clearing, which causes the associated version of $h_{i}$ to no longer exhibit point symmetry. In the latter case, interbank repayments must be factored into the value of liquid assets at clearing. This makes the corresponding version of $h_{i}$ dependent on the intricate interplay of liquidation strategies and payment requirements, thereby destroying its concave-convexity and precluding the application of the minimax theorem. Consequently, our primary research question is:
@@
Further notation.
All vectors are column vectors and denoted by bold symbols, e.g., ${\bf u}=(u_{1},\dots,u_{d})^{\top}\in\mathbb{R}^{d}$. We use $[m]$ and $\bf 1$} to represent the set $\{1,\dots,m\}$ and the vector of ones in a suitable dimension, respectively. For any two vectors ${\bf u},{\bf v}\in\mathbb{R}^{d}$, ${\bf u}\leq{\bf v}$ means an entry-wise inequality, ${\bf u},{\bf v}]=[u_{1},v_{1}]\times\cdots\times[u_{d},v_{d}]$, and ${\bf u}\wedge{\bf v}=(\min\{u_{1},v_{1}\},\dots,\min\{u_{d},v_{d}\})^{\top}$. For any matrix $\bf M$}, we denote its $i$-th column by ${\bf M}_{i}$. Finally, for any $x\in\mathbb{R}$, $x^{+}\coloneqq\max\{x,0\}$.
4 Robust Optimal Strategy with Permanent Price Impact

### 17 Mar 2026 v2

# Robust Optimal Strategies for Early Liquidation in Financial Systems

@@
Abstract
@@
Keywords: Systemic risk, Fire sales, Strategic liquidation, Worst-case approach, Permanent price impact, Interbank exposures
1.  Introduction
The global financial crisis in 2007-2008 provoked intense discussions on the factors that threatened the stability of the financial system. In the majority of relevant studies, two channels of risk contagion have been widely acknowledged as significant factors contributing to this instability. One such channel involves direct debt exposures among financial institutions. When an institution fails to meet debt obligations due to adverse economic shocks, its creditor institutions experience financial distress due to the resulting repayment shortfall (Eisenberg and Noe, 2001). Furthermore, during the crisis of asset erosion, financial institutions with inadequate liquid assets are forced to sell off substantial illiquid assets to satisfy their debt claims. Such liquidation overwhelms shallow market demand and triggers sharp declines in asset prices. This phenomenon, termed price impact, imposes significant losses on institutions holding the same illiquid assets, even without direct debt exposures among them. These fire sales constitute another channel of risk transmission and are recognized as a salient ingredient of the crisis (Clerc et al., 2016). Although the first channel has been studied more extensively in the literature, the second channel has gained increasing attention in recent years, particularly as financial institutions have actively reduced their debt exposures to other banks since the said crisis.
Our study departs from a conventional assumption in the existing literature that liquidation is conducted only at clearing. This assumption stems from the observation that liquidity-constrained institutions tend to favor borrowing over asset sales to avoid the prohibitive costs of liquidation. However, since borrowing capacity is not always guaranteed, relying solely on this option exposes banks to future funding shortages, particularly during systemic crises when credit conditions deteriorate (Acharya et al., 2011). In this case, preemptive liquidation prior to clearing becomes a rational solution. Then, this naturally raises a critical question: what volume of illiquid assets should an institution sell early to ensure its solvency? The answer is not immediately evident as liquidation outcomes are driven by the interplay of institutional liquidation decisions that jointly determine sale prices and are constrained by repayment requirements.
@@ 1.  Introduction
2.  Literature Review
@@
Impact of fire sales on systemic risk. Our research builds upon the literature on risk contagion via fire sales of illiquid assets, initiated by Cifuentes et al. (2005). They investigate the equilibrium of liquidation prices in financial systems where banks are forced to liquidate assets to satisfy regulatory constraints. Chen et al. (2016) formulate this equilibrium as an optimization problem and introduce a “liquidity amplifier” via sensitivity analysis, which quantifies how illiquidity intensifies fire-sale effects, thereby offering policy implications for intervention. In addition, Amini et al. (2016) establish sufficient conditions for the uniqueness of the equilibrium liquidation price, while Weber and Weske (2017) characterize the liquidation price in more complex settings where bankruptcy costs and cross-holdings are taken into account. These papers mainly focus on the assessment of systemic risk rather than the decision-making in liquidation and assume that liquidation takes place only at clearing. By contrast, we accommodate liquidation that occurs before clearing and offer novel insights into strategic liquidation decisions. A distinct line of work, pioneered by Greenwood et al. (2015), leverages correlations in banks’ asset exposures to measure fire-sale spillover effects. Subsequent studies have enriched this paradigm by modeling the dynamic process of fire sales (Duarte and Eisenbach, 2021) and incorporating partial information about asset holdings (Pang and Veraart, 2023). Although their approaches open up new directions, our modeling focus remains on extending the framework of Cifuentes et al. (2005), leaving the other models for further research.
@@
Strategic liquidation in financial systems. Our work contributes to the growing literature of strategic liquidation under systemic risk. Feinstein (2017) studies a financial system with multiple illiquid assets in which firms strategically choose asset sales to maximize their own valuations, and establishes the existence of equilibrium liquidation strategies. To analyze risk contagion arising from strategic liquidation, Braouezec and Wagalath (2019) consider a static game in which banks maximize their expected profits subject to regulatory capital requirements and characterize the resulting Nash equilibrium. Bichuch and Feinstein (2019) formulate an optimization problem in which borrowing and liquidation decisions are jointly determined to minimize total financing costs. While these studies analyze strategic liquidation through ex-post equilibrium outcomes of institutional interactions assuming rational behaviors of institutions, our work focuses on the ex-ante liquidation decision of an individual institution adopting a worst-case approach to the characterization of early liquidation strategies. [^1] It is essential to distinguish the notion of strategic liquidation considered here from the classical literature on optimal execution (e.g., Bertsimas and Lo, 1998; Gatheral and Schied, 2013), which focuses on minimizing trading costs when unwinding a fixed asset position over a given period under exogenous market illiquidity. In our systemic context, by contrast, illiquidity is not merely a frictional cost but an endogenous channel of risk contagion that amplifies price declines and threatens systemic stability.
@@
The role of permanent price impact in financial systems. Our research is closely related to studies on liquidation-induced systemic risk in the presence of permanent price impact, where price declines caused by liquidation do not immediately reverse and losses from liquidation accumulate over time. Cont and Schaanning (2019) incorporate such cumulative impacts to empirically measure indirect contagion from fire sales empirically. Capponi et al. (2020) examine a dynamic setting where persistent price impacts can trigger a self-reinforcing cycle of mutual fund redemptions and further liquidations. In the context of collateralized systems, Ghamami et al. (2022) and Pang and Veraart (2025) model clearing processes in which collateral is liquidated in multiple rounds, with price impacts from early sales carrying over entirely to subsequent rounds. While these works assume fully permanent price impact, implying that depressed prices do not recover, our work allows for partial price recovery. Furthermore, we examine robust liquidation decisions under permanent price impact from an individual decision-making perspective, whereas the aforementioned studies primarily focus on measuring systemic risk or elucidating contagion mechanisms.
@@
Contagion through interbank exposures. A large body of literature on systemic risk investigates the impact of interbank obligations. The seminal framework of Eisenberg and Noe (2001) models interbank repayments using a fixed-point characterization. Subsequent research has extended this framework along several dimensions while retaining its core analytical structure. For instance, Rogers and Veraart (2013) incorporate bankruptcy costs into the framework and explore incentives for consortium-based bailouts. Acemoglu et al. (2015) and Capponi et al. (2016) investigate how network topology and concentration affect shock transmission, respectively. Kusnetsov and Veraart (2019) and Veraart (2020) extend the framework to accommodate multiple maturities and quantify distress versus default contagion, respectively. Furthermore, this approach has been applied to diverse areas such as systemic stress scenario selection (Ahn et al., 2023), debt valuation (Barucca et al., 2020; Banerjee and Feinstein, 2022), intervention policies (Ahn and Kim, 2019; Bernard et al., 2022), and multilateral netting effects (Ahn, 2020; Amini et al., 2020). We also incorporate interbank exposures as an additional risk contagion channel. However, in contrast to the aforementioned works that analyze systemic risk from a regulatory perspective, our work examines the impact of this contagion channel on early liquidation decisions from the standpoint of an individual bank.
3.  Research Questions
@@
An illustrative example. Consider a system of $m$ banks operating over two periods ($t=1,2$), where liabilities are cleared at the end of the time horizon. Each bank $i$ possesses liquid assets valued at $e_{i}$ and holds $z_{i}$ units of illiquid assets. If the liquid assets prove insufficient to meet liabilities $L_{i}$ (i.e., $e_{i}<L_{i}$), the bank must cover the shortfall by selling off its illiquid holdings at a depressed price—a process known as fire sales. In contrast to the standard literature that assumes no early fire sales, we allow each bank $i$ to strategically liquidate its illiquid holdings $x_{i}$ at $t=1$, prior to clearing at $t=2$, in order to mitigate price impacts at clearing and sell a smaller amount of illiquid assets than in the scenario without early liquidation. Given a realization ${\bf x}=(x_{1},\ldots,x_{m})^{\top}$ of all banks’ early liquidation decisions, let $\psi_{i}({\bf x})$ denote the amount of illiquid assets bank $i$ must sell at clearing to fulfill its debt obligation for each $i=1,\ldots,m$. Then, by defining ${\bf x}_{-i}\coloneqq(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{m})^{\top}$, bank $i$’s liquid asset value at clearing becomes:
@@
Robust decision-making. In this example, each bank $i$ would seek to determine its strategy $x_{i}$ that optimizes $\phi_{i}(x_{i},{\bf x}_{-i})$ while maintaining robustness against other banks’ strategies ${\bf x}_{-i}$, as these are practically unobservable. However, as well documented in the literature (Amini et al., 2016) and as we shall demonstrate later, the mapping $\psi_{j}(\cdot)$ is characterized by a fixed-point equation involving nonconvex functions and is hence analytically intractable. A natural approach to this issue would be to replace $\psi_{j}({\bf x})$ by its surrogate $\psi_{j}({\bf 0})-x_{j}$ for all $j$, where $\bf 0$} denotes the zero vector. [^2] The early liquidation size $x_{j}$ is capped at $\psi_{j}({\bf 0})$ for all $j=1,\ldots,m$. This is because, as noted earlier, the goal of early liquidation is to reduce the amount of illiquid assets sold and the losses from liquidation compared to the case without it. Similarly, $\psi_{j}({\bf x})$ is capped at $\psi_{j}({\bf 0})-x_{j}$. This accounts for the worst-case scenario in which banks eventually sell off the same amount of illiquid assets as they would in the absence of early liquidation. Then, the said decision-making problem of bank $i$ can be formulated as follows:
@@
Solving (2) is surprisingly simple. By letting $s_{i}=\sum_{j\neq i}x_{j}$ and $\bar{s}_{i}=\sum_{j\neq i}\bar{x}_{j}$, it can be recast as $\max_{x_{i}\in[0,\bar{x}_{i}]}\min_{s_{i}\in[0,\bar{s}_{i}]}g_{i}(x_{i},s_{i})$, where $g_{i}(x,s)\coloneqq e_{i}+xQ(x+s)+(\bar{x}_{i}-x)Q(\bar{x}_{i}+\bar{s}_{i}-(x+s_{i}))$. Under the common assumptions on $Q(\cdot)$ in the literature (see Assumption 1), $g_{i}(x,s)$ is concave in $x$ and convex in $s$. Therefore, the minimax theorem (Rockafellar, 2015, Theorem 36.3) holds: $\max_{x_{i}\in[0,\bar{x}_{i}]}\min_{s_{i}\in[0,\bar{s}_{i}]}g_{i}(x_{i},s_{i})=\min_{s_{i}\in[0,\bar{s}_{i}]}\max_{x_{i}\in[0,\bar{x}_{i}]}g_{i}(x_{i},s_{i})$. Furthermore, $h_{i}$ satisfies the following point symmetry: $g_{i}(x_{i},s_{i})=g_{i}(\bar{x}_{i}-x_{i},\bar{s}_{i}-s_{i})$ for any $x_{i}\in[0,\bar{x}_{i}]$} and $s_{i}\in[0,\bar{s}_{i}]$}. Accordingly, $(\bar{x}_{i}/2,\bar{s}_{i}/2)$ is the saddle point of $h_{i}$, implying that
@@
Main issues and questions. The simplicity of identifying the robust optimal strategy, however, is limited to the special case presented above. When practical considerations are incorporated, finding the strategy becomes significantly more complex. Specifically, this paper focuses on two distinct extensions prevalent in the literature: incorporating permanent price impact and including interbank exposures. In the former case, early liquidation exerts downward pressure on the asset price at clearing, which causes the associated version of $h_{i}$ to no longer exhibit point symmetry. In the latter case, interbank repayments must be factored into the value of liquid assets at clearing. This makes the corresponding version of $h_{i}$ dependent on the intricate interplay of liquidation strategies and payment requirements, thereby destroying its concave-convexity and precluding the application of the minimax theorem. Consequently, our primary research question is:
@@
Further notation. All vectors are column vectors and denoted by bold symbols, e.g., ${\bf u}=(u_{1},\dots,u_{d})^{\top}\in\mathbb{R}^{d}$. We use $[m]$ and $\bf 1$} to represent the set $\{1,\dots,m\}$ and the vector of ones in a suitable dimension, respectively. For any two vectors ${\bf u},{\bf v}\in\mathbb{R}^{d}$, ${\bf u}\leq{\bf v}$ means an entry-wise inequality, ${\bf u},{\bf v}]=[u_{1},v_{1}]\times\cdots\times[u_{d},v_{d}]$, and ${\bf u}\wedge{\bf v}=(\min\{u_{1},v_{1}\},\dots,\min\{u_{d},v_{d}\})^{\top}$. For any matrix $\bf M$}, we denote its $i$-th column by ${\bf M}_{i}$. Finally, for any $x\in\mathbb{R}$, $x^{+}\coloneqq\max\{x,0\}$.
4.  Robust Optimal Strategy with Permanent Price Impact